### v1

# Generalized de Branges-Rovnyak spaces

Abstract
Given the reproducing kernel $k$ of the Hilbert space ${{\mathcal{H}}_{k}}$ we study spaces ${{\mathcal{H}}_{k}}(b)$ whose reproducing kernel has the form $k(1-bb^{*})$, where $b$ is a row-contraction on ${{\mathcal{H}}_{k}}$. In terms of reproducing kernels this it the most far-reaching generalization of the classical de Branges-Rovnyaks spaces, as well as their very recent generalization to several variables. This includes the so called sub-Bergman spaces [zhu_subbergm1] in one or several variables. We study some general properties of ${{\mathcal{H}}_{k}}(b)$ e.g. when the inclusion map into ${{\mathcal{H}}_{k}}$ is compact. Our main result provides a model for ${{\mathcal{H}}_{k}}(b)$ reminiscent of the Sz.-Nagy-Foiaş model for contractions (see also [aleman_conbackshift]). As an application we obtain sufficient conditions for the containment and density of the linear span of $\{k_{w}:w\in{\mathcal{X}}\}$ in ${{\mathcal{H}}_{k}}(b)$. In the standard cases this reduces to containment and density of polynomials. These methods resolve a very recent conjecture [gu_hypercontractions] regarding polynomial approximation in spaces with kernel $\frac{(1-b(z)b(w)^{*})^{m}}{(1-z\overline{w})^{\beta}},1\leq m<\beta,m\in{\mathbb{N}}$.
1 Introduction
Motivated by models for contractions, de Branges and Rovnyak initiated in the 60’s [dbr65, dbr66] the study of the Hilbert spaces with a reproducing kernel of the form
(1) $k(z,w)=\frac{1-b(z)b(w)^{*}}{1-z\overline{w}},\quad z,w\in{\mathbb{D}},$
Here, as usual, ${\mathbb{D}}$ denotes the unit disc in the complex plane and the operator-valued function $b:{\mathbb{D}}\to{\mathcal{B}}({\mathbb{C}},l^{2})$ is analytic with operator norm bounded by one in ${\mathbb{D}}$. These ideas have been considerably developed by Sarason [sarason] who studied extensively these spaces in the case when $b$ is a scalar-valued analytic function, and called them sub-Hardy spaces ${\mathcal{H}}(b)$. For example, he represented them as ranges of the defect operators $(I-T_{b}T_{b}^{*})^{1/2}$ endowed with the norm making these operators co-isometric. Here $T_{b}$ denotes the (analytic) Toeplitz operator with symbol $b$ on the Hardy space $H^{2}$.
${\mathcal{H}}(b)$-spaces attracted a lot of attention during the years, a good account on those developments can be found in the book by Fricain and Mashreghi [fricain-mashreghi] and in the references therein. An important idea of Sarason [sarason] reveals a kind of dichotomy regarding the structure of these spaces. Roughly speaking, it turns out that the properties of ${\mathcal{H}}(b)$ are either similar to those of a shift invariant subspace of $H^{2}$, or to those of the orthogonal complement of such a subspace (model space). The first alternative occurs when $b$ is not an extreme point of the unit ball of $H^{\infty}$, the algebra of bounded analytic functions in ${\mathbb{D}}$, while the second occurs when $b$ is an extreme point of that set. In fact, it is proved in [sarason] that the following are equivalent:
  1. (i)
    $b\in{\mathcal{H}}(b)$,
  2. (ii)
    ${\mathcal{H}}(b)$ is invariant under the shift operator defined by $Sf(z)=zf(z)$,
  3. (iii)
    the polynomials are contained in ${\mathcal{H}}(b)$ and are dense in the space.
We shall refer to the above situation as Sarason’s dichotomy.
Zhu [zhu_subbergm1], [zhu_subbergm2] developed these ideas by considering Toeplitz operators $T_{b}$ acting on the Bergman space, i.e. the range of $(I-T_{b}T_{b}^{*})^{1/2}$ with the appropriate norm. We should point out that these ranges are quite different. For example, it turns out that $\operatorname{Ran}(I-T_{b}T_{b}^{*})=\operatorname{Ran}(I-T_{b}^{*}T_{b})$, which fails dramatically in the Hardy space context.
Such spaces are called sub-Bergman spaces and by analogy to (1) they have the reproducing kernel
(2) $k(z,w)=\frac{1-b(z)\overline{b(w)}}{(1-z\overline{w})^{2}},$
i.e. it is obtained by replacing in (1) the Szegő kernel by the Bergman kernel. All of these spaces are invariant for the forward shift and contain the polynomials. For our purposes it is important to note that in [chu_density] Chu proved the density of polynomials in all sub-Bergman spaces.
This idea generated a lot of further work [zhu_subbergm2, zhu_subbergm3, abkar-jafarzadeh, symesak] mainly devoted to extensions to weighted Bergman spaces in one or several variables.
Another important and very recent development regards the generalization of the original de Branges-Rovnyak spaces to the context of several complex variables. From an operator-theoretic point of view, the most natural generalization of the Szegő kernel to the unit ball in several complex variables is the Drury-Arveson kernel
$k(z,w)=\frac{1}{1-\langle z,w\rangle},\quad z,w\in{\mathbb{B}}_{d}.$
Jury and Martin considered in [jury-martin_extremal] reproducing kernel Hilbert spaces with kernel of the form
(3) $k_{b}(z,w)=\frac{1-b(z)b(w)^{*}}{1-\langle z,w\rangle},$
where $b$ is a row-vector pointwise multiplier of the Drury-Arveson space. They characterized the multipliers $b$ for which the equivalent assertions (i), or (ii) from above hold, showing in particuar, that the Sarason dichotomy does occur in this case. Shortly after, Hartz [hartz_columnrow] completed their work by showing that the equivalent statements (i) and (ii) above hold precisely when $b$ is an extreme point of the unit ball of the multiplier space. However, the density of polynomials (assertion (iii)) when (i) or (ii) hold, remains open in this context.
Spaces with kernel of the form (3) appear also in the recent doctoral thesis of Sautel as models for certain commuting tuples which are expansive on the complement of their null-space. Moreover, together with Richter [sautel] they provide sufficient conditions under which a kernel of the form (3) is a complete Nevanlinna-Pick kernel.
The present paper is concerned with a very general construction of this type. More precisely, we consider a reproducing kernel Hilbert space with reproducing kernel $k$ on a non-void set ${\mathcal{X}}$, denoted by ${\mathcal{H}}_{k}$. In order to deal with vector-valued $b$’s as well, we assume that $b$ is a non-constant contractive multiplier of ${\mathcal{H}}_{k}\otimes\ell^{2}$ into ${\mathcal{H}}_{k}$. This implies (see §2.1 below) that
$k^{b}(x,y)=(1-b(x)b(y)^{*})k(x,y),\quad x,y\in{\mathcal{X}},$
is the reproducing kernel of a Hilbert function space on ${\mathcal{X}}$ which we denote by ${\mathcal{H}}_{k}(b)$. It could be called sub-${\mathcal{H}}_{k}$ space.
Adopting this general point of view reveals that a number of properties of the particular spaces mentioned above, especially in the sub-weighted Bergman case, only depend on general results about reproducing kernels. For example, in §3.2 we consider the embedding of ${\mathcal{H}}_{k}(b)$ into ${\mathcal{H}}_{k}$ which is always contractive. We prove that a necessary condition for the compactness of this embedding is that $\|b(x_{n})\|\to 1$, whenever $k(x_{n},x_{n})\to\infty$.
In the case when ${\mathcal{H}}_{k}$ is a space of analytic functions on the unit disc ${\mathbb{D}}$ with $k(z_{n},z_{n})\to\infty$ whenever $|z_{n}|\to 1$ and $b$ is a scalar-valued multiplier, the above condition implies that $b$ is a finite Blaschke product. In Theorem 4 we show that this condition is also sufficient provided that it is sufficient in the simplest case $b(z)=z$. This is a fairly far-reaching generalization of results of Zhu [zhu_subbergm1, zhu_subbergm3] proved for standard weighted Bergman spaces.
The results extend also to the context of several complex variables, namely when $k$ is a power of the Drury-Arveson kernel and $b$ is an automorphism of the unit ball. We also supply some examples showing that, in general, the above condition fails to be sufficient.
Beyond extending these interesting results, the study of general ${\mathcal{H}}_{k}(b)$- spaces has additional motivations. In this paper we shall focus on the following two aspects:
  1. 1)
    Reproducing kernel structure induced by operator inequalities,
  2. 2)
    Reproducing kernels with a complete Nevanlinna-Pick factor.
1) When $b(0)=0$ the original de Branges-Rovnyak spaces ${\mathcal{H}}(b)$ with $b$ scalar-valued are precisely the spaces of analytic functions on ${\mathbb{D}}$ with reproducing kernel normalized at 00 where the backward shift $L$ is contractive and, in addition, $I-LL^{*}$ has rank one. The recent paper [aleman_conbackshift] provides a thorough study of such spaces without the last assumption, which yields Hilbert spaces with reproducing kernel of the general form (1). Moreover, it turns out that the case when $\operatorname{rank}(I-LL^{*})<\infty$ leads to a theory analogous to the one of sub-Hardy spaces, in particular Sarason’s dichotomy holds. Extending these ideas to the context of several variables is obviously very complicated due to the intricate structure of left inverses to the row operator given by multiplication of the coordinates. For this reason, the expansivity of that row operator plays the central role in Sautel’s work [sautel] mentioned above.
The second hereditary inequality is more involved and was derived by Shimorin in [shimorin_wold]. If $S$ denote the forward shift on a space of analytic functions ${\mathcal{H}}$ on the disc with normalized reproducing kernel at the origin, Shimorin’s condition reads
(4) $\|Sf+g\|^{2}\leq 2(\|f\|^{2}+\|Sg\|^{2}).$
It turns out (see also §3.1 below) that it characterizes general sub-Bergman spaces obtained by replacing the scalar function $b$ in (2) by an operator-valued one. In §3.1 we prove a generalization of both inequalities. Assume that ${\mathcal{H}}$ is a reproducing kernel Hilbert space of analytic functions in ${\mathbb{D}}$ with normalized kernel at the origin. We show that the inequality
$\left\|f_{0}+\sum_{n\geq 1}u_{n}f_{n}\right\|^{2}\leq\|\varphi f_{0}\|^{2}+\sum_{n\geq 1}\|f_{n}\|^{2},\quad f_{n}\in{\mathcal{H}},\,n\geq 0,$
where $\varphi,u_{n},~{}n\geq 1$ are fixed multipliers, characterizes ${\mathcal{H}}_{k}(b)$-spaces, with $k$ a kernel of Bergman-type, a class introduced by McCullough and Richter in [MR].
2) By a normalized complete Nevanlinna-Pick (CNP) kernel on the nonvoid set ${\mathcal{X}}$, normalized at $z_{0}\in{\mathcal{X}}$ we mean a kernel of the form
$(z,w)\mapsto\frac{1}{1-b(z)b(w)^{*}},\qquad b(z)b(w)^{*}=\langle b(z),b(w)\rangle_{\ell^{2}},\quad z,w\in{\mathcal{X}},$
where $\|b(z)\|_{l^{2}}\leq 1,~{}z\in{\mathcal{X}},~{}b(z_{0})=0$. Such reproducing kernels arise from solving certain interpolation problems for multipliers [agler-mccarthy]. These kernels have a number of remarkable properties and have attracted a lot of attention recently. Here we are interested in a much larger class, namely those which have a normalized CNP factor, i.e. the reproducing kernels $k$ for which there is a normalized CNP kernel $s$ such that $k/s$ is a reproducing kernel as well (positive definite). Such kernels play a crucial for the very general commutant lifting theorem by S. Shimorin [shimorin_leech] and very recently in [aleman_cnpfactor] it has been shown that functions in spaces with such kernels can be represented as quotients of multipliers. Given $k$ with the CNP factor $\frac{1}{1-b(z)b(w)^{*}}$ the ”quotient” kernel
$k_{b}(z,w)=k(z,w)(1-b(z)b(w)^{*}),$
coresponds to the ${\mathcal{H}}_{k}(b)$ space which actually appears in Shimorin’s commutant lifting theorem as well. One of the main difficulties in dealing with such spaces is to understand which functions belong to them other than the reproducing kernels. In §4.1 we give sufficient conditions such that ${\mathcal{H}}_{k}(b)$ contains the kernels $k_{y}\in{\mathcal{H}}_{k}$ or even more, such that the span of these kernels is dense in ${\mathcal{H}}_{k}(b)$ (Theorem 6). The proof of this approximation theorem is based on an analogue of the Sz.-Nagy-Foiaş model inspired by the work in [aleman_conbackshift]. The idea is interesting in its own right and is presented in §3.3.
There is a large class of examples satisfying these conditions, like the sub-Hardy spaces ${\mathcal{H}}(b)$ with $b$ non-extremal in the unit ball of $H^{\infty}$ and all ${\mathcal{H}}_{k}(b)$ spaces with ${\mathcal{H}}_{k}$ a weighted Bergman space on a domain in ${\mathbb{C}}^{d}$ and $b$ a non-constant analytic row contraction on that domain. Intuitively speaking, the kernels $k_{y}\in{\mathcal{H}}_{k},~{}y\in{\mathcal{X}}$ play the same role as polynomials in the theory of spaces of analytic functions on the $d$-dimensional unit ball. In particular, these results extend Chu’s density theorem [chu_density] to sub-Bergman spaces on very general domains in ${\mathbb{C}}^{d}$. In fact, for the standard weighted Bergman spaces on the ball we show in §5.1 that Sarason’s dichotomy does not occur, that is conditions (i), (ii) and (iii) hold in any of such sub-weighted Bergman spaces.
Finally, these results and ideas apply to an interesting class of kernels considered recently in [gu_hypercontractions], namely positive integer powers of the sub-weighted Bergman kernels
$\frac{(1-b(z)b(w)^{*})^{m}}{(1-z\overline{w})^{\beta}}\qquad m\in{\mathbb{N}},m<\beta\in{\mathbb{R}}.$
These kernels are naturally connected to $m$-hypercontractive operators on standard weighted Bergman spaces. In [gu_hypercontractions] it was conjectured that polynomials are dense in the corresponding spaces of analytic functions on ${\mathbb{D}}$. As an application of the results mentioned above we prove that the conjecture holds true. Our methods yield a proof in the higher dimensional case as well.
The paper is organised as follows. Section 2 serves to a preliminary purpose. Section 3 contains general properties of ${{\mathcal{H}}_{k}}(b)$, in §3.1 we discuss the operator inequalities mentioned before. §3.2 is devoted to the embedding of ${{\mathcal{H}}_{k}}(b)$ in ${{\mathcal{H}}_{k}}$ and the cases where it is compact. The analogue of the Sz.-Nagy-Foiaş model is described in §3.3. Section 4 contains the general approximation results, while in Section 5 we apply these ideas to general weighted Bergman spaces. The proof of the conjecture in [gu_hypercontractions] is deferred to section 6.
2 Preliminaries
2.1 Reproducing kernels
For the sake of completeness we recall some basic properties of these objects. The material is standard, can be found in [aronszajn] and [paulsen]. If ${\mathcal{X}}$ denotes a non-empty set, a function $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{C}}$ is called positive (write $k\gg 0$) if for all finite subsets $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$ the matrix $[k(x_{i},x_{j})]_{i,j=1}^{n}$ is positive semi-definite. Given $k\gg 0$ there exists a unique Hilbert space ${{\mathcal{H}}_{k}}$ consisting of complex-valued functions on ${\mathcal{X}}$, such that point-evaluations are bounded linear functionals on ${{\mathcal{H}}_{k}}$ given by
$f(y)=\langle f,k_{y}\rangle_{{{\mathcal{H}}_{k}}},\quad x,y\in{\mathcal{X}},$
where we have used the common notation $k_{y}=k(\cdot,y)$. For this reason, $k$ is called a reproducing kernel.
It is easy to verify that if $k,h$ are reproducing kernels on ${\mathcal{X}}$ with $k-h\gg 0$, then ${\mathcal{H}}_{h}$ is contractively contained in ${{\mathcal{H}}_{k}}$. In fact, the map $k_{y}\to h_{y}$ extends by linearity to a contraction, whose adjoint is the inclusion map. This yields the following result.
Proposition 1.
Let $k$ be a reproducing kernel on ${\mathcal{X}}$, then $f:{\mathcal{X}}\to{\mathbb{C}}$ is in ${\mathcal{H}}_{k}$ if and only if
$c^{2}k(x,y)-f(x)\overline{f(y)}\gg 0$
for some $c>0$ and the least such $c$ is equal to $\|f\|_{{\mathcal{H}}_{k}}$.
The classical Schur’s product theorem for positive semi-definite matrices applies to reproducing kernels and gives the following result.
Theorem 1 (Schur’s product theorem).
Let $k$ and $l$ be reproducing kernels on ${\mathcal{X}}$. Then $kl$ is also a reproducing kernel.
Definition 1.
Let $k$ be a reproducing kernel on ${\mathcal{X}}$. A function $\varphi:{\mathcal{X}}\to{\mathbb{C}}$ is a multiplier of ${\mathcal{H}}_{k}$ if $\varphi f\in{\mathcal{H}}_{k}$ for every $f\in{\mathcal{H}}_{k}$. The algebra of these operators is denoted $\operatorname{Mult}({{\mathcal{H}}_{k}})$.
Note that if $\varphi\in\operatorname{Mult}({{\mathcal{H}}_{k}})$, then the induced linear map $M_{\varphi}$ on ${\mathcal{H}}_{k},~{}M_{\varphi}f=\varphi f$ is bounded by the closed graph theorem. Moreover, we have
(5) $M_{\varphi}^{*}k_{y}=\overline{\varphi(y)}k_{y}.$
Consequently, $\sup_{x\in{\mathcal{X}}}|\varphi(x)|\leq\|M_{\varphi}\|$. Using also the fact that the linear span of $\{k_{y}:y\in{\mathcal{X}}\}$ is dense in ${{\mathcal{H}}_{k}}$ one can derive the following result.
Proposition 2.
Let $k$ be a reproducing kernel on ${\mathcal{X}}$, then $\varphi$ is a multiplier of ${\mathcal{H}}_{k}$ if and only if
$k(x,y)(c^{2}-\varphi(x)\overline{\varphi(y)})\gg 0$
for some $c>0$ and the least such $c$ is equal to $\|M_{\varphi}\|$.
Given two sets ${\mathcal{F}},{\mathcal{G}}$ of complex-valued functions on ${\mathcal{X}}$ shall consider the linear span of their pointwise products, that is,
$\operatorname{span}{\mathcal{F}}\cdot{\mathcal{G}}=\operatorname{span}\{fg:f\in{\mathcal{F}},g\in{\mathcal{G}}\}.$
The next result is essentially a consequence of the above. We include a short proof for the convenience of the reader.
Proposition 3.
Let $s,t$ be reproducing kernels on ${\mathcal{X}}$ and $k=st$. Then:
  1. (i)
    $\operatorname{span}\operatorname{Mult}({\mathcal{H}}_{s})\cdot\operatorname{Mult}({\mathcal{H}}_{t})\subset\operatorname{Mult}({\mathcal{H}}_{k})$, in particular $\operatorname{Mult}({\mathcal{H}}_{s})\subset\operatorname{Mult}({\mathcal{H}}_{k})$ contractively,
  2. (ii)
    If $f\in{\mathcal{H}}_{s}$ and $g\in{\mathcal{H}}_{t}$, then $fg\in{\mathcal{H}}_{k}$ and $\|fg\|_{{\mathcal{H}}_{k}}\leq\|f\|_{{\mathcal{H}}_{s}}\|g\|_{{\mathcal{H}}_{t}}$, in particular if the constant function $1\in{\mathcal{H}}_{t}$, then ${\mathcal{H}}_{s}$ is boundedly contained in ${\mathcal{H}}_{k}$,
  3. (iii)
    If ${\mathcal{F}}$ and ${\mathcal{G}}$ are dense sets in ${\mathcal{H}}_{s}$ and ${\mathcal{H}}_{t}$, respectively, then $\operatorname{span}{\mathcal{F}}\cdot{\mathcal{G}}$ is dense in ${\mathcal{H}}_{k}$.
Proof.
(i) Let $\varphi\in\operatorname{Mult}({\mathcal{H}}_{s})$ and $\psi\in\operatorname{Mult}({\mathcal{H}}_{t})$, then by Proposition 2
$s(x,y)\|M_{\varphi}\|^{2}\gg s(x,y)\varphi(x)\overline{\varphi(y)}$ $\gg 0,$
$t(x,y)\|M_{\psi}\|^{2}\gg t(x,y)\psi(x)\overline{\psi(y)}$ $\gg 0.$
By Schur’s product theorem, the order is preserved when multiplying the left-hand side and the right-hand side, respectively. That is
$k(x,y)\|M_{\varphi}\|^{2}\|M_{\psi}\|^{2}\gg\varphi(x)\psi(x)\overline{\varphi(y)\psi(y)}$
and it follows that $\varphi\psi\in\operatorname{Mult}({\mathcal{H}}_{k})$. Clearly, any element in $\operatorname{span}\operatorname{Mult}({\mathcal{H}}_{s})\cdot\operatorname{Mult}({\mathcal{H}}_{t})$ is a sum of products as above. The contractive containment follows by letting above $\psi=1$ which gives
$k(x,y)(\|M_{\varphi}\|^{2}-\varphi(x)\overline{\varphi(y)})\gg 0.$
(ii) is very similar and will be omitted.
(iii) By (ii) we have that ${\mathcal{F}}\cdot{\mathcal{G}}\subset{\mathcal{H}}_{k}$. If $y\in{\mathcal{X}},~{}\varepsilon>0$, and $f\in{\mathcal{F}},~{}g\in{\mathcal{G}}$ with
$\|f-s_{y}\|_{{\mathcal{H}}_{s}}<\varepsilon,\quad\|g-t_{y}\|_{{\mathcal{H}}_{t}}<\varepsilon,$
by the first part of (ii) we have
$\|fg-k_{y}\|_{{\mathcal{H}}_{k}}$ $\leq\|g\|_{{\mathcal{H}}_{t}}\|f-s_{y}\|_{{\mathcal{H}}_{s}}$
$\leq(\varepsilon+\|t_{y}\|_{{\mathcal{H}}_{t}})\varepsilon+(\varepsilon+\|s_{y}\|_{{\mathcal{H}}_{s}})\varepsilon$
which gives the desired result by the density of the linear span of the kernel functions $k_{y},\,y\in{\mathcal{X}}$ in ${\mathcal{H}}_{K}$. ∎
A repeated direct application of the last part of the proposition leads to the following result.
Corollary 1.
Let ${\mathcal{F}}$ be an algebra of functions on ${\mathcal{X}}$ and let $s^{j},~{}j=1,\ldots,n$ be reproducing kernels on ${\mathcal{X}}$ such that ${\mathcal{F}}$ is dense in ${\mathcal{H}}_{s^{j}},j=1,\ldots,n$. If $k=\prod_{j=1}^{n}s^{j}$ then ${\mathcal{F}}$ is dense in ${\mathcal{H}}_{k}$.
We close the paragraph with a few words about the vector-valued case. The reproducing kernel of a Hilbert spaces of vector-valued functions is an operator valued function of two variables with similar properties to the scalar-valued version.
In this paper we shall be concerned with a very special case. Given a scalar reproducing kernel $k$ on ${\mathcal{X}}$, we shall consider the Hilbert space ${{\mathcal{H}}_{k}}\otimes\ell^{2}$, that is, the space of functions $f:{\mathcal{X}}\to\ell^{2}$ such that
$f=\begin{pmatrix}f_{0}\\ f_{1}\\ \vdots\end{pmatrix},\,f_{j}\in{{\mathcal{H}}_{k}},j\geq 0,\qquad\|f\|_{{{\mathcal{H}}_{k}}\otimes\ell^{2}}^{2}=\sum_{j\geq 0}\|f_{j}\|_{{\mathcal{H}}_{k}}^{2}<\infty.$
This Hilbert space has the operator-valued reproducing kernel $(x,y)\mapsto I_{\ell^{2}}k(x,y)$. In other words, given $f\in{{\mathcal{H}}_{k}}\otimes\ell^{2},~{}y\in{\mathcal{X}},~{}e\in\ell^{2}$, we have
$\langle f(y),e\rangle_{\ell^{2}}=\langle f,k_{y}e\rangle_{{{\mathcal{H}}_{k}}\otimes\ell^{2}}.$
A multiplier $\varphi$ from ${{\mathcal{H}}_{k}}\otimes\ell^{2}$ to ${{\mathcal{H}}_{k}}$ is row-vector valued function such that $\varphi f\in{{\mathcal{H}}_{k}}$ for every $f\in{{\mathcal{H}}_{k}}\otimes\ell^{2}$. We write $\varphi\in\operatorname{Mult}({{\mathcal{H}}_{k}}\otimes\ell^{2},{{\mathcal{H}}_{k}})$. In this case (5) becomes
(6) $M_{\varphi}^{*}k_{y}=\varphi(y)^{*}k_{y}.$
Note that $\varphi(y)^{*}$ can be interpreted as a column-vector. Finally, let us also mention that Proposition 2 extends verbatim to this context by simply replacing $\overline{\varphi(y)}$ by $\varphi(y)^{*}$.
2.2 Some unitarily invariant kernels on the unit ball of ${\mathbb{C}}^{d}$
Let $d$ be a positive integer and
${\mathbb{B}}_{d}=\{z\in{\mathbb{C}}^{d}:\|z\|<1\},$
be open the unit ball in ${\mathbb{C}}^{d}$. A kernel $k$ on ${\mathbb{B}}_{d}$ is invariant under the unitary group on ${\mathbb{C}}^{d}$ if and only if $k(z,w)$ is a function of $\langle z,w\rangle.$
An important class of examples, see [aleman_besov], are reproducing kernels of radially weighted Besov spaces
[…]
Here $R^{s}$ is a power of the radial derivative operator $R=\sum_{j=1}^{d}z_{j}\frac{\partial}{\partial z_{j}}$, $V$ is the Lebesgue measure on the ball, and $\omega$ is a radial integrable function on ${\mathbb{B}}_{d}$. It is easy to verify that monomials form an orthogonal basis in these spaces. A particular case is the Drury-Arveson kernel on ${{\mathbb{B}}_{d}}$ is defined by
$s(z,w)\coloneq\frac{1}{1-\langle z,w\rangle},\qquad z,w\in{{\mathbb{B}}_{d}}.$
When $d=1$ it is the reproducing kernel for the Hardy space $H^{2}$ on the unit disk ${\mathbb{D}}$ and $s$ is actually called the Szegő kernel. For $d>1$, the corresponding space is called the Drury-Arveson space and plays a key role in the theory of commuting operator tuples. The recent survey [hartz_invitation] contains a lot of interesting material in this direction.
For any $\beta>0$ observe that
$s(z,w)^{\beta}=\sum_{n=0}^{\infty}a_{n}\langle z,w\rangle^{n},$
where $a_{n}\geq 0$ therefore by Schur’s product theorem $s^{\beta}$ is a reproducing kernel.
Definition 2.
Let $\beta>0$ we write ${{\mathcal{H}}_{\beta}}$ for the space of holomorphic functions on ${{\mathbb{B}}_{d}}$ that corresponds to the reproducing kernel
$s^{\beta}(z,w)\coloneq\frac{1}{(1-\langle z,w\rangle)^{\beta}}$
These spaces are extensively studied in the literature. If $\beta>d$, then ${\mathcal{H}}_{s^{\beta}}$ is a (standard) weighted Bergman space on ${{\mathbb{B}}_{d}}$ as noted in the next paragraph.
2.3 Weighted Bergman spaces on a domain of ${\mathbb{C}}^{d}$
Let $\Omega$ be a domain in ${\mathbb{C}}^{d}$ and let $\mu$ be a regular Borel positive measure on $\Omega$ such that the analytic functions in $L^{2}(\mu)$ form a closed subspace where point-evaluations are continuous linear functionals. These spaces are usually called weighted Bergman spaces and we denote them by $L^{2}_{a}(\mu)$. By assumption the space $L^{2}_{a}(\mu)$ has a reproducing kernel which we denote by $k^{\mu}$, and, as pointed out before, the norm is given by
[…]
If $\beta>d$, the spaces ${{\mathcal{H}}_{\beta}}$ from the previous paragraph satisfy ${{\mathcal{H}}_{\beta}}=L^{2}_{a}(\mu)$ with $\Omega={\mathbb{B}}_{d}$, $d\mu=c_{\alpha}(1-|z|^{2})^{\alpha}dV$, where $\alpha=\beta-(d+1)$, and $c_{\alpha}$ is a normalizing constant.
The vector valued version $L^{2}_{a}(\mu)\otimes\ell^{2}$ is defined with help of the norm
[…]
One verifies that $\operatorname{Mult}(L^{2}_{a}(\mu))$ equals the algebra of bounded analytic functions on $\Omega$ with equality of norms. Similarly, $\operatorname{Mult}(L^{2}_{a}(\mu)\otimes\ell^{2},L^{2}_{a}(\mu))$ consists of bounded analytic row-vectors and the multiplier norm coincides with the supremum norm. For our purposes it is important to note that these multipliers never attain their norm unless the symbols are constant. Indeed, if $\|bf\|=\|b\|_{\infty}\|f\|$ for some non-zero $f\in L^{2}_{a}(\mu)\otimes\ell^{2}$, then $\|b\|_{\infty}=\|b(z)\|_{\ell^{2}}$ for all $z\in\Omega$, which easily implies that $b$ is constant. In the terminology of [aleman_freeouter] these Bergman spaces have no non-trivial sub-inner multipliers. This is in strong contrast to smaller reproducing kernel spaces on the ball. In fact, more than that is true under a very mild condition on $\mu$. If $L_{a}^{2}(\mu)$ is a reproducing kernel closed subspace of $L^{2}(\mu)$ then there exists a measure $\tilde{\mu}$ such that $L_{a}^{2}(\tilde{\mu})\otimes\ell^{2}=L_{a}^{2}(\mu)\otimes\ell^{2}$ and for each compact subset $K$ of $\Omega$ there exists $c(K)>0$ with
(7) $\tilde{\mu}|K\geq c(K)V|K,$
where $V$ denotes the Lebesgue measure on ${\mathbb{C}}^{d}$. For example, one can choose a positive $v\in L^{1}(V)$ with essinf $\textup{essinf }V|K>0$ for any compact subset $K$ of $\Omega$ and set
$d\tilde{\mu}=d\mu+\frac{v(z)}{1+\|k^{\mu}_{z}\|^{2}}dV.$
Proposition 4.
Let $b:\Omega\to\ell^{2}$ be analytic nonconstant and such that $b(z)$ is a row-contraction for every $z\in\Omega$. If $\mu$ satisfies (7), then for every $z\in\Omega$ there exists $c_{z}>0$ such that
$\|f\|^{2}-\|bf\|^{2}\geq c_{z}\|f(z)\|_{\ell^{2}}^{2},$
for all $f\in L^{2}_{a}(\mu)\otimes\ell^{2}$.
Proof.
Let $z\in\Omega$ and $B_{z}$ be an open ball in ${\mathbb{C}}^{d}$ centered at $z$ with $\overline{B_{z}}\subset\Omega$. By the assumtion on $b$ there is $a_{z}>0$ such that $1-\|b(\zeta)\|_{\ell^{2}}^{2}\geq a_{z}$, $\zeta\in B_{z}$. This implies for each $e\in\ell^{2}$,
$\|e\|_{\ell_{2}}^{2}-|b(\zeta)e|^{2}\geq(1-\|b(\zeta)\|_{\ell^{2}}^{2})\|e\|_{\ell^{2}}^{2}\geq a_{z}\|e\|_{\ell^{2}}^{2}.$
Using also (7) we have for $f\in L^{2}_{a}(\mu)\otimes\ell^{2}$
[…]
$\leq\frac{\|f\|^{2}-\|bf\|^{2}}{V(B_{z})c(\overline{B_{z}})a_{z}},$
which completes the proof. ∎
3 Basic properties of generalized de Branges-Rovnyak spaces
Let us recall the general notion of a de Branges-Rovnyak space pointed out in the Introduction. Given be a scalar-valued reproducing kernel $k$ on the non-empty set ${\mathcal{X}}$ let ${\mathcal{H}}_{k}$ denote the corresponding Hilbert function space.
Definition 3.
Let $b\in\operatorname{Mult}({\mathcal{H}}_{k}\otimes\ell^{2},{\mathcal{H}}_{k})$ with $\|M_{b}\|\leq 1$. The sub-${\mathcal{H}}_{k}$ space, or the de Branges-Rovnyak space with respect to ${\mathcal{H}}_{k}$ and $b$ is the Hilbert space corresponding to the reproducing kernel
$k^{b}(x,y)\coloneq k(x,y)(1-b(x)b(y)^{*}),\qquad b(x)b(y)^{*}=\langle b(x),b(y)\rangle_{\ell^{2}},$
which is positive due to the multiplier criterion. We denote this space by ${\mathcal{H}}_{k}(b)$.
One can verify that as a set, ${\mathcal{H}}_{k}(b)$ is equal to $\operatorname{Ran}(I-M_{b}M_{b}^{*})^{\frac{1}{2}}$ and its norm satisfies
$\langle(I-M_{b}M_{b}^{*})^{\frac{1}{2}}f,(I-M_{b}M_{b}^{*})^{\frac{1}{2}}g\rangle_{{\mathcal{H}}_{k}(b)}=\langle f,g\rangle_{{\mathcal{H}}_{k}},$
whenever $f,g\in\operatorname{Ran}(I-M_{b}M_{b}^{*})^{\frac{1}{2}}={\mathcal{H}}_{k}\ominus\ker(I-M_{b}M_{b}^{*})^{\frac{1}{2}}.$ Furthermore, ${\mathcal{H}}_{k}(b)$ is contractively contained in ${\mathcal{H}}_{k}$, this can be seen either from the above formula, or the fact that
$k(z,w)-k^{b}(z,w)=b(x)b(y)^{*}k(x,y),$
which is positive by Schur’s product theorem.
3.1 Examples via operator inequalities
It is well known that the backward shift
$Lf(z)=\frac{f(z)-z}{z},$
is a contraction on any de Branges-Rovnyak space ${\mathcal{H}}(b)={\mathcal{H}}_{s}(b)$, with $s(z,w)=\frac{1}{1-\overline{w}z}$ and $b(0)=0$. The starting point of the work in [aleman_conbackshift] is the observation that the converse holds true as well, as the following result shows.
Proposition 5 ([aleman_conbackshift]).
Let ${\mathcal{H}}$ be a Hilbert space of analytic functions on ${\mathbb{D}}$ with a reproducing kernel $k$ normalized at 00. $L$ is contractive on ${\mathcal{H}}$ if and only if there exists an analytic row-contraction $b$ with $b(0)=0$ such that
$k(z,w)=\frac{1-b(z)b(w)^{*}}{1-z\overline{w}}.$
For reproducing kernel Hilbert spaces of analytic functions in ${\mathbb{D}}$ having the identity function $z$ as a multiplier and a reproducing kernel normalized at the origin, the contractivity of $L$ is equivalent to the fact that the forward shift $M_{z}$ is expansive, that is, $\|zf\|\geq\|f\|$. In particular, such spaces have a general de Branges-Rovnyak kernel as above.
There is a very recent extension of this idea to the context of several complex variables due to S. Richter and J. Sautel [sautel]. It uses the Drury-Arveson kernel $s(z,w)=\frac{1}{1-\langle z,w\rangle}$ instead of the Szegö kernel. The proof is much more involved in this case.
Theorem 2.
Let $T=(T_{1},\ldots,T_{d}):{\mathcal{H}}^{d}\to{\mathcal{H}}$ be a bounded row operator on an infinite dimensional Hilbert space ${\mathcal{H}}$, and let $m\in{\mathbb{N}}$. T is unitarily equivalent to $(M_{z},{\mathcal{H}}_{s}(b)$ for some $b\in\operatorname{Mult}({\mathcal{H}}_{s}\otimes\mathcal{D},{\mathcal{H}}_{s}\otimes\mathcal{E})$ and Hilbert spaces $\mathcal{D}$ and $\mathcal{E}$ such that $b(0)=0$, $\dim\mathcal{E}=m$, and that $M_{z}$ is a bounded row operator if and only if $T$ satisfies the following four conditions:
  1. (i)
    $T_{i}T_{j}=T_{j}T_{i}$ for all $1\leq i,j\leq d$,
  2. (ii)
    $\dim([\operatorname{Ran}(T-\lambda)]^{\perp})=m$ for every $\lambda\in{{\mathbb{B}}_{d}}$,
  3. (iii)
    $\|Tx\|_{\mathcal{H}}\geq\|x\|_{{\mathcal{H}}^{d}}$ for all $x\in\operatorname{Ker}(T)^{\perp}$,
  4. (iv)
    $\bigcap\limits_{n\geq 0}\sum\limits_{|w|=n}T_{w}{\mathcal{H}}=\{0\}$.
Sub-Bergman spaces are also related to an operator inequality which is considerably more subtle. It is due to S. Shimorin [shimorin_wold] and it asserts that on the (unweighted) Bergman space on ${\mathbb{D}}$, the forward shift $M_{z}f=zf$ satisfies
(8) $\|zf+g\|^{2}\leq 2(\|f\|^{2}+\|zg\|^{2}).$
As the previous inequalities, (8) is hereditary in the sense that it continues to hold for restrictions of the operator involved to its invariant subspaces. This estimate has far-reaching consequences regarding invariant subspaces of the Bergman shift. A crucial fact (see also [AHR]) is that (8) is closely related to sub-Bergman kernels in the following way.
Proposition 6 (Shimorin’s inequality ).
Let ${\mathcal{H}}={{\mathcal{H}}_{k}}$ be a Hilbert space of analytic functions on ${\mathbb{D}}$ such that the reproducing kernel $k$ is normalized at 00, $M_{z}$ and $L$ are bounded on ${\mathcal{H}}$. Then the inequality (8) holds on ${\mathcal{H}}$ if and only if there exists an analytic row-contraction $b$ with $b(0)=0$ such that
(9) $k(z,w)=\frac{1-b(z)b(w)^{*}}{(1-z\overline{w})^{2}},$
In order to study $M_{z}$-invariant subspaces of the Bergman space one needs to drop both assumptions in the above proposition, that is, to consider non-normalized kernels, as well as left inverses of the forward shift with a null space of arbitrary dimension.
Several interesting extensions of Shimorin’s inequality have been found by Olofsson and Wenman (see [OW] and the references therein).
A very general class of kernels containing all the above examples are the so- called Bergman-type kernels introduced by McCullough and Richter in [MR] These are reproducing kernels in ${\mathbb{D}}$ of the form
(10) $k(z,w)=\frac{1}{1-\varphi(z)\overline{\varphi(w)}(1-u(z)u(w)^{*})},\quad z,w\in{\mathbb{D}},$
with $\varphi(0)=0$ and $k(z,z)\to\infty$ when $|z|\to 1^{-}$. It is shown in [MR] that with $\rho(z)=\frac{\varphi(z)}{z}$ can be chosen to be outer and bounded from above and below. Clearly, the Szegö kernel is obtained for $\varphi(z)=z,u=0$, and the unweighted Bergman kernel for $\varphi(z)=z\sqrt{2},u(z)=\frac{z}{\sqrt{2}}$.
Also, if $k$ is of Bergman-type we have $Mult({\mathcal{H}}(k))=H^{\infty}$ with equality of norms (see again [MR]). The following theorem extends all previous results in one complex variable to the general context of Bergman-type kernels.
Theorem 3.
Let $k$ be a Bergman-type kernel on ${\mathbb{D}}$ of the form (10). Let ${\mathcal{H}}$ be a Hilbert space of analytic functions on ${\mathbb{D}}$ whose reproducing kernel is normalized at 00 and assume that $\varphi\in Mult({\mathcal{H}})$, the range of $M_{\varphi}$ is closed of codimension $1$, and that $u\in Mult({\mathcal{H}}\otimes\ell^{2})$. Then ${\mathcal{H}}={\mathcal{H}}_{k}(b)$ for an analytic row-contraction $b$ in ${\mathbb{D}}$ with $b(0)=0$, if and only
(11) $\left\|f_{0}+\sum_{n\geq 1}u_{n}f_{n}\right\|^{2}\leq\|\varphi f_{0}\|^{2}+\sum_{n\geq 1}\|f_{n}\|^{2},$
whenever $f_{0},f_{1}\ldots,f_{n},\ldots\in{\mathcal{H}}$.
Proof.
The assumptions on $M_{\varphi}$ imply that its range equals the space of functions in ${\mathcal{H}}$ vanishing at the origin. Consequently, the operator $L_{\varphi}$ defined by
$L_{\varphi}f=\frac{f-f(0)}{\varphi},\quad f\in{\mathcal{H}},$
is a bounded left-inverse of $M_{\varphi}$. Then the inequality (11) is equivalent to
(12) $\left\|L_{\varphi}f_{0}+\sum_{n\geq 1}u_{n}f_{n}\right\|^{2}\leq\|f_{0}\|^{2}+\sum_{n\geq 1}\|f_{n}\|^{2}.$
Indeed, if (11) holds, we apply it to $L_{\varphi}f_{0},f_{1},\ldots$. From the fact that the kernel of ${\mathcal{H}}$ is normalized we obtain
$\|\varphi L_{\varphi}f_{0}\|^{2}=\|f_{0}-f_{0}(0)\|^{2}\|\leq\|f_{0}\|,$
and (12) follows.
Conversely, if (12) holds we apply it to $\varphi f_{0},f_{1},\ldots$ and obtain (11). Now (12) is further equivalent to the fact that the operator $T:\bigoplus_{\mathbb{N}}{\mathcal{H}}\to{\mathcal{H}}$ defined by
$T\sum_{n\geq 0}f_{n}=L_{\varphi}f_{0}+\sum_{n\geq 1}u_{n}f_{n},$
is contractive. This can be re-written as
$I-TT^{*}\geq 0\,\Leftrightarrow(I-TT^{*})h\gg 0\,\Leftrightarrow(I-TT^{*})h=bb^{*},$
where $h$ is the reproducing kernel in ${\mathcal{H}}$ and $b$ is an analytic row-contraction in ${\mathbb{D}}$. Since
$L_{\varphi}L_{\varphi}^{*}h(z,w)=\frac{h(z,w)-1}{z\overline{w}},\quad M_{u}M_{u}^{*}h(z,w)=u(z)\overline{u(w)}h(z,w),\,u\in Mult({\mathcal{H}}),$
we can solve for $h$ in $(I-TT^{*})h=bb^{*}$ and obtain after a straightforward calculation
$h(z,w)=(1-b(z)b^{*}(w))k(z,w),$
i.e. $h$ is the reproducing kernel of ${\mathcal{H}}_{k}(b)$. ∎
Besides the Szegő, or the unweighted Bergman kernel, some standard examples of Bergman-type kernels are
[…]
These examples are discussed in [OW]. The corresponding Hilbert spaces are standard weighted Bergman spaces. Of course, there are many other examples in this class, in particular in [MR] there are produced examples such that the forward shift on the corresponding space is not subnormal.
3.2 Compact embedding
As mentioned in the beginning of this section ${\mathcal{H}}_{k}(b)$ is always contained contractively in ${\mathcal{H}}_{k}$. It is natural to ask when the contractive inclusion map is compact. As pointed out in the Introduction this question attracted attention in the context of sub-weighted Bergman spaces.
If the original space ${\mathcal{H}}_{k}$ is smaller, then the question is less interesting as we shall see below. We start with a simple necessary condition for the compactness of the inclusion map which holds under mild assumptions in the most general context of a kernel $k$ on the nonvoid set ${\mathcal{X}}$.
Proposition 7.
Let $k$ be a reproducing kernel on a non-empty set ${\mathcal{X}}$ with the property that $k_{y}$ is a bounded function for every $y\in{\mathcal{X}}$. Let $b\in\operatorname{Mult}({\mathcal{H}}_{k}\otimes\ell^{2})$ and suppose that the inclusion map from ${\mathcal{H}}_{k}(b)$ to ${\mathcal{H}}_{k}$ is compact. Then for every sequence $(x_{n})$ in ${\mathcal{X}}$ with $\lim_{n\to\infty}k(x_{n},x:n)=\infty$, we have
$\lim_{n\to\infty}\|b(x_{n})\|_{\ell^{2}}=1.$
Proof.
Let us begin with the simple observation that if $J:{\mathcal{H}}_{k}(b)\to{\mathcal{H}}_{k}$ is the inclusion map,then its adjoint $J^{*}$ satisfies
$j^{*}k_{y}(x)=(1-b(x)b^{*}(y))k_{y}(x),\quad x,y\in{\mathcal{X}}.$
Now let $(x_{n})$ be a sequence as in the statement and set $f_{n}=\frac{k_{x_{n}}}{\|k_{x_{n}}\|}$, so that $\|f_{n}\|=1$. Furthermore, by the assumption on $k$ we have
$\langle f_{n},k_{y}\rangle=\frac{k(x_{n},y)}{\|k_{x_{n}}\|}=\frac{\overline{k_{y}(x_{n})}}{\|k_{x_{n}}\|}\to 0\text{as $n\to\infty$}.$
This proves that $(f_{n})$ tends weakly to 00 in ${\mathcal{H}}_{k}$, hence if $J$ is compact, so is $J^{*}$, that is, $J^{*}f_{n}$ must tend in norm to 0. Thus,
$\|J^{*}f_{n}\|^{2}=1-\|b(x_{n})\|^{2}\to 0.$
and the result follows. ∎
The result is of particular interest for most common reproducing kernel Hilbert spaces of analytic functions in the unit disc, in the case when $b$ is scalar-valued. Indeed, if ${\mathcal{H}}_{k}$ is such a space with
$k(z,z)\to\infty,\quad|z|\to 1^{-},$
and $b\in Mult({\mathcal{H}}_{k})$ is contractive, by Proposition 7 the compactness of the inclusion $J:{\mathcal{H}}_{k}(b)\to{\mathcal{H}}_{k}$ implies that the bounded analytic function $b$ satisfies for all $\lambda\in\mathbb{T}$
$\lim_{z\to\lambda}|b(z)|=1,$
hence $b$ is a finite Blaschke product. However, in this generality, finite Blaschke products may fail to be contractive multipliers so the corresponding generalized de Branges-Rovnyak spaces might not exist.
Example 1.
  1. a)
    In the Dirichlet space $D$ with kernel
    $-\frac{1}{z\overline{w}}\log(1-z\overline{w}),\quad z,w\in{\mathbb{D}},$
    every finite Blaschke product $b$ is a multiplier, but (see [RS])
    $\|bf\|>\|f\|,\quad f\in D\setminus\{0\}.$
  2. b)
    Let ${\mathcal{H}}_{2}$ be the unweighted Bergman space and ${\mathcal{H}}_{1}=H^{2}$ be the Hardy space on the unit disc. Consider the reproducing kernel Hilbert space ${\mathcal{H}}$ consisting of analytic functions $f$ in ${\mathbb{D}}$ which can be written (uniquely) as
    $f(z)=u_{1}(z^{2})+zu_{2}(z^{2}),\quad u_{1}\in{\mathcal{H}}_{1},\,u_{2}\in{\mathcal{H}}_{2},$
    with norm
    $\|f\|^{2}=\|u_{1}\|^{2}_{{\mathcal{H}}_{1}}+\|u_{2}\|^{2}_{{\mathcal{H}}_{2}}.$
    Then obviously, $M_{z}$ is unbounded on ${\mathcal{H}}$ while if $\varphi\in H^{\infty}$, the function $\psi(z)=\varphi(z^{2})$ is a multiplier with $\|M_{\psi}\|=\|\psi\|_{\infty}.$
Proposition 8.
Given a reproducing kernel Hilbert space ${\mathcal{H}}_{k}$ of analytic functions in ${\mathbb{D}}$, the finite Blaschke products are contractive multipliers if and only if the identity function $\zeta(z)=z$ is a contractive multiplier.
Proof.
One direction is obvious since the identity is a finite Blaschke product, while the converse follows by standard functional calculus for contractions from the equality $b(M_{z})=M_{b}$ valid for any finite Blaschke product. ∎
A similar result can be derived for the compactness of the embedding of ${\mathcal{H}}_{k}(b)$ into ${\mathcal{H}}_{k}$, but the proof is somewhat more involved. Our next result provides a far-reaching generalization of the result in [zhu_subbergm3].
Theorem 4.
Assume that ${\mathcal{H}}_{k}$ is a reproducing kernel Hilbert space of analytic functions on ${\mathbb{D}}$ such that $M_{z}$ is contractive on it. Then if $b\in Mult({\mathcal{H}}_{k})$ is contractive, the inclusion map From ${\mathcal{H}}_{k}(b)$ into ${\mathcal{H}}_{k}$ is compact if and only if $I-M_{z}M_{z}^{*}$ is compact.
The proof is essentially based on the identity below. A similar formula appears in [gu_hypercontractions]. Given $a\in{\mathbb{D}}$ we denote by $\varphi_{a}$ the disc-automorphism
$\varphi_{a}(z)=\frac{z-a}{1-\overline{a}z},\quad z\in{\mathbb{D}}.$
Lemma 1.
Let $s(z,w)=\frac{1}{1-z\overline{w}}$ and let $b$ be a finite Blaschke product with zeros $a_{1},\ldots,a_{n}\in{\mathbb{D}}$. Then
(13) $(1-b(z)\overline{b(w)})s(z,w)=\sum_{l=1}^{n}(\varphi_{a_{l}}^{\prime}(z)\overline{\varphi_{a_{l}}^{\prime}(w)})^{\frac{1}{2}}\prod_{j=1}^{l-1}\varphi_{a_{j}}(z)\overline{\varphi_{a_{j}}(w)}.$
Proof.
We proceed by induction. For $a\in{\mathbb{D}}$ a direct computation gives
$(1-\varphi_{a}(z)\overline{\varphi_{a}(w)})s(z,w)=(\varphi_{a}^{\prime}(z)\overline{\varphi_{a}^{\prime}(w)})^{\frac{1}{2}},$
which is the case $n=1$. Suppose that (13) holds for any Blaschke product with $n$ zeros, and let $b$ be the finite Blaschke product with zeros $a_{1},\ldots,a_{n+1}\in{\mathbb{D}}$. Let $b_{0}=1$ and let $b_{k},~{}k\geq 1,$ denote the Blaschke product having only the first $k$ zeros, $a_{1},\ldots,a_{k}$. Write
$1-b(z)\overline{b(w)}=1-b_{n}(z)\overline{b_{n}(w)}+b_{n}(z)\overline{b_{n}(w)}(1-\varphi_{a_{n+1}}(z)\overline{\varphi_{a_{n+1}}(w)}).$
By the induction hypotheses
$(1-b_{n}(z)\overline{b_{n}(w)})s(z,w)=\sum_{j=1}^{n}(\varphi_{a_{l}}^{\prime}(z)\overline{\varphi_{a_{l}}^{\prime}(w)})^{\frac{1}{2}}b_{l-1}(z)\overline{b_{l-1}(w)},$
and by the case $n=1$,
$b_{n}(z)\overline{b_{n}(w)}(1-\varphi_{a_{n+1}}(z)\overline{\varphi_{a_{n+1}}(w)})s(z,w)=(\varphi_{a_{n+1}}^{\prime}(z)\overline{\varphi_{a_{n+1}}^{\prime}(w)})^{\frac{1}{2}}b_{n}(z)\overline{b_{n}(w)},$
which proves the induction step and the result follows. ∎
Proof of Theorem 4. The assumption $\|M_{z}\|\leq 1$ implies that for all $a\in{\mathbb{D}}$, the functions $u_{a}(z)=(1-\overline{a}z)^{-1},~{}z\in{\mathbb{D}}$ are multipliers, in fact $M_{u_{a}}=(I-\overline{a}M_{z})^{-1}$. Assume that $I-M_{z}M_{z}^{*}$ is compact and let $b$ be a finite Blaschke product with zeros $a_{1},\ldots,a_{n}\in{\mathbb{D}}$. Multiply both sides of (13) by $k(z,w)/s(z,w)$, use the identity $u(z)\overline{u(w)}k(z,w)=M_{u}M_{u}^{*}k_{w}(z)$ together with the density of the linear span of $\{k_{w}:~{}w\in{\mathbb{D}}\}$ to obtain an equality of the form
(14) $I-M_{b}M_{b}^{*}=\sum_{l=1}^{n}M_{v_{l}}(I-M_{z}M_{z}^{*})M_{v_{l}}^{*},$
with $v_{1},\ldots,v_{n}\in Mult({\mathcal{H}}_{k})$. Thus, by assumption, each operator in the sum is compact, hence, so is $I-M_{b}M_{b}^{*}$. Now if $J:{\mathcal{H}}_{k}(b)\to{\mathcal{H}}_{k}$ is the inclusion map, as above it follows that
$J^{*}k_{w}(z)=(1-b(z)\overline{b(w)})k(z,w),$
hence $JJ^{*}=I-M_{b}M_{b}^{*}$ is compact. Conversely, if $JJ^{*}=I-M_{b}M_{b}^{*}$ is compact, by Proposition 7, $b$ must be a Blaschke product, so that (14) holds. Note that all operators on the right hand side are positive, and in the first term we have
$v_{1}(z)=(\varphi_{a_{1}}^{\prime}(z))^{\frac{1}{2}}=\frac{(1-|a_{1}|^{2})^{\frac{1}{2}}}{1-\overline{a_{1}}z},$
is an invertible multiplier. This leads to
$M_{v_{1}}^{-1}(I-M_{b}M_{b}^{*})(M_{v_{1}}^{-1})^{*}\geq I-M_{z}M_{z}^{*}.$
Since the left hand side is compact and the right hand side is positive it follows that $I-M_{z}M_{z}^{*}$ is compact and the proof is complete. ∎
The theorem applies not only to arbitrary weighted Bergman spaces on ${\mathbb{D}}$, or the Hardy space. The result holds in many other reproducing kernel spaces, for example those corresponding to radial kernels
$k(z,w)=\sum_{n\geq 0}k_{n}(z\overline{w})^{n},\quad z,w\in{\mathbb{D}},$
with $10<k_{n}\leq k_{n+1},~{}\lim_{n\to\infty}\frac{k_{n}}{k_{n+1}}=$ 1 .
It is an interesting question to characterize the compact embedding of ${\mathcal{H}}_{k}(b)$ into ${\mathcal{H}}_{k}$ for a a contractive row multiplier $b$.
The situation is much more complicated for analytic reproducing kernels on the unit ball ${\mathbb{B}}_{d},~{}d>1$. A particular case that can be treated with the method above is when $b$ is an automorphism of ${{\mathbb{B}}_{d}}$ and ${\mathcal{H}}_{k}={\mathcal{H}}_{\beta}$ for some $\beta>1$. In this case we have the following identity similar to (13) (see for example [rudin_uball]),
$1-\langle\psi(z),\psi(w)\rangle=\frac{(1-\langle a,a\rangle)(1-\langle z,w\rangle)}{(1-\langle z,a\rangle)(1-\langle a,w\rangle)}$
where $a=b^{-1}(0)$. This leads to
$s^{\beta}(z,w)(1-\psi(z)\psi(w)^{*})=s^{\beta-1}(z,w)\varphi(z)\overline{\varphi(w)},$
where $\varphi(z)=\frac{(1-\langle a,a\rangle)^{\frac{1}{2}}}{1-\langle z,a\rangle}$ is an invertible multiplier of ${\mathcal{H}}_{\beta-1}$. In other words, ${\mathcal{H}}_{\beta}(b)={\mathcal{H}}_{\beta-1}$ with equivalent norms. This proves the following result.
Proposition 9.
If $\beta>1$ and $b$ is an automorphism of ${{\mathbb{B}}_{d}}$, then ${{\mathcal{H}}_{\beta}}(b)$ is compactly embedded in ${{\mathcal{H}}_{\beta}}$.
3.3 A model for ${{\mathcal{H}}_{k}}(b)$
One of the main difficulies that occur when dealing with de-Branges-Rovnyak spaces and their generalizations is finding functions in the space other than finite linear combinations of reproducing kernels. In some cases this led to a number of recent non-trivial results. For example, the main result in [aleman_conbackshift] shows that in one variable, ${\mathcal{H}}_{s}(b)$ contains a dense subset of functions that extend continuously to $\overline{{\mathbb{D}}}$. Further refinements of that result have been recently obtained by Limani and Malman [LM]. the method in [aleman_conbackshift] relies on a special formula for the norm in ${\mathcal{H}}_{s}(b)$ which turns out to be essentially related to the famous Nagy-Foiaş model for the backward shift on these spaces (see [NFBK]). Some of these ideas do extend to the general context of ${\mathcal{H}}_{k}(b)$-spaces and provide useful information about their structure. To illustrate this, consider a scalar-valued reproducing kernel $k$ on the non-empty set ${\mathcal{X}}$ with corresponding Hilbert space ${\mathcal{H}}_{k}$ and fix a contractive row-multiplier $b\in\operatorname{Mult}({\mathcal{H}}_{k}\otimes\ell^{2},{\mathcal{H}}_{k})$. The considerations below apply to a scalar, or finite dimensional vector $b$ is a trivial way. Therefore for the rest of the subsection we shall consider the case when $b\in\operatorname{Mult}({\mathcal{H}}_{k}\otimes\ell^{2},{\mathcal{H}}_{k})$.
Denote by $\Delta=(I-M_{b}^{*}M_{b})^{\frac{1}{2}}:{\mathcal{H}}_{k}\otimes\ell^{2}\to{\mathcal{H}}_{k}\otimes\ell^{2}$. If $\Delta$ (or $I-M^{*}_{b}M_{b}$) is injective, so that
$\|f\|_{\Delta}=\|\Delta f\|,$
is a norm on ${\mathcal{H}}_{k}\otimes\ell^{2}$. In this case we let ${\mathcal{L}}_{\Delta}$ be the completion of ${\mathcal{H}}_{k}\otimes\ell^{2}$ with respect to $\|\cdot\|_{\Delta}$. We shall be interested in the closed subspace ${\mathcal{M}}\subset{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ defined by
(15) =${\mathcal{M}}=\{(bf,f):f\in{\mathcal{H}}_{k}\otimes\ell^{2}\}$
Let us record two simple facts about this subspace.
Proposition 10.
Suppose that $\Delta$ is injective and let ${\mathcal{M}}\subset{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ be the subspace given by (15). Then:
  1. (i)
    ${\mathcal{M}}$ is closed,
  2. (ii)
    For $(u,v)\in{\mathcal{H}}_{k}\oplus{\mathcal{H}}_{k}\otimes\ell^{2}$ its orthogonal projection onto ${\mathcal{M}}$, $P_{\mathcal{M}}(u,v)$ is given by
(16) $P_{\mathcal{M}}(u,v)=(bw,w),\qquad w=M_{b}^{*}u+\Delta^{2}v.$
Proof.
(i) If $((bf_{n},f_{n}))$ converges in ${\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ then from
$\|(bf_{n},f_{n})-(bf_{m},f_{m})\|^{2}=\|b(f_{n}-f_{m})\|^{2}_{{\mathcal{H}}_{k}}+\|\Delta(f_{n}-f_{m})\|^{2}_{{\mathcal{H}}_{k}\otimes\ell^{2}}=\|f_{n}-f_{m}\|^{2}_{{\mathcal{H}}_{k}\otimes\ell^{2}},$
we see that $(f_{n})$ is Cauchy in ${\mathcal{H}}_{k}\otimes\ell^{2}$, hence convergent.
(ii) With $f\in{\mathcal{H}}_{k}\otimes\ell^{2}$ and $w$ as in the statement we have
$\langle(bf,f),(bw,w)\rangle_{{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}}$ $=\langle f,w\rangle_{{\mathcal{H}}_{k}\otimes\ell^{2}}$
$=\langle f,M_{b}^{*}u+\Delta^{2}v\rangle_{{\mathcal{H}}_{k}\otimes\ell^{2}}$
$=\langle(bf,f),(u,v)\rangle{{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}}.$
The main result of this subsection is given below.
Theorem 5.
Suppose that $\Delta$ is injective and let ${\mathcal{M}}\subset{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ be the subspace given by (15). Then the map $J:{\mathcal{M}}^{\perp}\to{\mathcal{H}}_{k}(b)$, $J(u,v)=u$ is unitary. Equivalently, for every $f\in{\mathcal{H}}_{k}(b)$ there exists a unique $g\in{\mathcal{L}}_{\Delta}$ such that $(f,g)\in{\mathcal{M}}^{\perp}$ and
$\|f\|_{{\mathcal{H}}_{k}(b)}^{2}=\|f\|_{{\mathcal{H}}_{k}}^{2}+\|g\|_{{\mathcal{L}}_{\Delta}}^{2}.$
Proof.
Let us note first that if $(0,g)\in{\mathcal{M}}^{\perp},~{}g\in{\mathcal{L}}_{\Delta}$, then $g=0$. Indeed, in this case $g$ is orthogonal in ${\mathcal{L}}\Delta$ to the dense subspace ${\mathcal{H}}_{k}\otimes\ell^{2}$, hence $g=0$. In particular, if $P_{{\mathcal{M}}^{\perp}}=I-P_{\mathcal{M}}$, this implies that the linear span of $\{P_{{\mathcal{M}}^{\perp}}(k_{y},0):~{}y\in{\mathcal{X}}\}$ is dense in ${\mathcal{M}}^{\perp}$. By Proposition 10 (ii) we can easily calculate the elements of this set. Indeed, according to (16) we have
$P_{\mathcal{M}}(k_{y},0)=(bw,w),\qquad w=M_{b}^{*}k_{y},$
so that
$(I-P_{\mathcal{M}})(k_{y},0)=P_{{\mathcal{M}}^{\perp}}(k_{y},0)=(k_{y}-b(x)b^{*}(y)k_{y},-b^{*}(y)k_{y})=(k_{y}^{b},-b^{*}(y)k_{y}),$
where $k^{b}$ denotes the reproducing kernel in ${\mathcal{H}}_{k}(b)$. But then
$\langle P_{{\mathcal{M}}^{\perp}}(k_{y},0),P_{{\mathcal{M}}^{\perp}}(k_{z},0)\rangle_{{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\delta}}$
$=\langle k_{y}-bb^{*}(y)k_{y},k_{z}-bb^{*}(z)k_{z}\rangle_{{\mathcal{H}}_{k}}+\langle\Delta b^{*}(y)k_{y},b^{*}(z)k_{z}\rangle_{{\mathcal{H}}_{k}\otimes\ell^{2}}$
$=k_{y}(z)-b(z)b^{*}(y)k_{y}(z)-\overline{b(y)b^{*}(z)k_{z}(y)}+\langle bb^{*}(y)k_{y},bb^{*}(z)k_{z}\rangle_{{\mathcal{H}}_{k}}$
$\quad+b(z)b^{*}(y)k_{y}(z)-\langle bb^{*}(y)k_{y},bb^{*}(z)k_{z}\rangle_{{\mathcal{H}}_{k}}$
$=k_{y}(z)-b(z)b^{*}(y)k_{y}(z)=k_{y}^{b}(z).$
Thus the map $J$ in the statement preserves the norm on a dense subset of ${\mathcal{M}}^{\perp}$ and maps it onto a dense subset of ${\mathcal{H}}_{k}(b)$. The result follows.
The hypotheses that $\Delta$ is injective might seem restrictive. It does not apply to inner functions $b$ in the case of the Hardy space on ${\mathbb{D}}$, or, more generally to sub-inner multipliers of ${\mathcal{H}}_{k}$. It turns out that smaller spaces than $H^{2}$ on ${\mathbb{D}}$, like for example standard weighted Dirichlet spaces possess a large set of sub-inner multipliers (see [aleman_freeouter], Theorem 14.9). On the other hand, recall that weighted Bergman spaces do not have such multipliers.
4 Approximation results
4.1 The general approximation theorem
Given a space ${\mathcal{H}}_{k}(b)$, one of the main difficulties in understanding its structure is the lack of knowledge about its elements other than the reproducing kernels. Therefore it is natural to relate this problem to elements in the larger (”known”) space ${\mathcal{H}}_{k}$, in particular to the original kernels $k_{y},~{}y\in X$. We shall consider the map $J$ from Theorem 5, more precisely, its inverse $J^{-1}:{\mathcal{H}}_{k}(b)\to{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ and denote, as in the previous subsection, by ${\mathcal{M}}$ the closed subspace of ${\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}$ with
=${\mathcal{M}}=\{(bg,g):~{}g\in{\mathcal{H}}_{k}\otimes\ell^{2}\}.$
Proposition 11.
Assume that $\Delta$ is injective. If $y\in{\mathcal{X}}$ then $k_{y}\in{\mathcal{H}}_{k}(b)$ if and only if there exists $c_{y}>0$ such that
(17) $|b(y)g(y)|\leq c_{y}\|g\|_{\Delta},\quad g\in{\mathcal{H}}_{k}\otimes\ell^{2}.$
In this case, there exists a unique $l_{y}\in{\mathcal{L}}_{\Delta}$ with
(18) $\langle g,l_{y}\rangle_{{\mathcal{L}}_{\Delta}}=-b(y)g(y),\quad g\in{\mathcal{H}}_{k}\otimes\ell^{2},$
and $J^{-1}k_{y}=(k_{y},l_{y})$.
Proof.
If $k_{y}\in{\mathcal{H}}_{k}(b)$ and $J^{-1}k_{y}=(k_{y},l_{y})\in{\mathcal{M}}^{\perp}$, then $l_{y}$ must satisfy (18) which implies (17). Conversely, if (17) holds, the existence and uniqueness of $l_{y}$ follows by the Riesz representation theorem and Theorem 5 gived $k_{y}\in{\mathcal{H}}(b)$. ∎
Under the assumption $k_{y}\in{\mathcal{H}}(b),~{}y\in{\mathcal{X}}$, the natural question regarding approximation in this space is whether the linear span of these functions is dense. We provide a sufficient condition for this conclusion.
Theorem 6.
Suppose that $\Delta$ is injective and that:
  1. (i)
    If $y\in{\mathcal{X}},~{}e\in\ell^{2}$ there exists $d_{y}>0$ such that
(19) $|\langle g(y),e\rangle_{\ell^{2}}|\leq d_{y}\|g\|_{\Delta}\|e\|_{2},\quad g\in{\mathcal{H}}_{k}\otimes\ell^{2}.$
Equivalently, point evaluations are continuous on ${\mathcal{L}}_{\Delta}$.
  1. (ii)
    If $f\in{\mathcal{L}}_{\Delta}$ and $bf\in{\mathcal{H}}_{k}$ then $f\in{\mathcal{H}}_{k}\otimes\ell^{2}.$
Then the linear span of $\{k_{y}:y\in{\mathcal{X}}\}$ is contained and dense in ${\mathcal{H}}_{k}(b)$.
Proof.
The hypothesis (i) together with Proposition 11 imply that $k_{y}\in{\mathcal{H}}_{k}(b),\,y\in{\mathcal{X}}$. Now suppose that $f\in{\mathcal{H}}_{k}(b)$ annihilates $k_{y},~{}y\in{\mathcal{X}}$. Let
$J^{-1}f=(f,g),\,\,J^{-1}k_{y}=(k_{y},l_{y})\in{\mathcal{M}}^{\perp},\quad y\in{\mathcal{X}},$
and apply Proposition 11 to conclude that (18) holds for all $y\in{\mathcal{X}}$. Thus
$0=\langle f,k_{y}\rangle_{{\mathcal{H}}_{k}(b)}=\langle J^{-1}f,J^{-1}k_{y}\rangle_{{\mathcal{H}}_{k}\oplus{\mathcal{L}}_{\Delta}}=f(y)-b(y)g(y),$
for all $y\in{\mathcal{X}}$, i.e. $f=bg\in{\mathcal{H}}_{k}$. But then by assumption (ii), we have $g\in{\mathcal{H}}_{k}\otimes\ell^{2}$ and $J^{-1}f=(f,g)=(bg,g)\in{\mathcal{M}}$ which implies $(f,g)=0$ and the result follows. ∎
Remark 1.
  1. 1)
    In the classical case when ${\mathcal{H}}_{k}$ equals the Hardy space $H^{2}$ and $b$ is a scalar analytic function bounded by $1$ in ${\mathbb{D}}$, assumption (i) is equivalent to the fact that $b$ is not extremal in $H^{\infty}$. This follows immediately by the standard Szegö theorem. Moreover, in this case there exists an outer function $a$ with $|a|^{2}+|b|^{2}=1$ a.e. on the unit circle, which easily implies that (ii) holds automatically.
  2. 2)
    In the general case, an analogue of the above condition is to assume that there exists a multiplier $a$ such that $I-M_{b}^{*}M_{b}\geq M_{a}^{*}M_{a}$ (see for example [jury-martin_extremal] and the discussion in the Introduction). If $a$ is zero-free then this assumption implies assumption (i) in the theorem above. However, we do not know whether it implies (ii) as well.
5 Weighted sub-Bergman spaces
Let $\Omega$ be a domain of ${\mathbb{C}}^{d}$, $\mu$ be a positive measure on $\Omega$ such that the Bergman space $L^{2}_{a}(\mu)$ is a closed subspace of $L^{2}(\mu)$ with reproducing kernel $k^{\mu}$ (see section 2.3. For an analytic row-contraction $b$ on $\Omega\subset{\mathbb{C}}^{d}$ we consider the sub- Bergman space ${\mathcal{H}}_{k^{\mu}}(b)$. The aim of this subsection is to show that the general result given in Theorem 6 does always apply to these spaces.
Theorem 7.
Let $b$ be any analytic row-contraction on $\Omega$ with $\|b(z_{0})\|<1$ for some $z_{0}\in\Omega$. Then the linear span of $\{k^{\mu}_{z}:~{}z\in\Omega\}$ is dense in ${\mathcal{H}}_{k^{\mu}}(b)$.
Proof.
The statement is obvious when $b$ is a constant with $\|b\|<1$, so we may assume $b$ is not constant. We want to apply Theorem 6. To this end we need to verify the suumptions (i) and (ii) in that theorem. Assumption (i) follows by an application of Proposition 4. To verify (ii), let $g\in{\mathcal{L}}_{\Delta}$ with $bg\in L_{a}^{2}(\mu)$ and let $(g_{n})$ be a sequence in $L_{a}^{2}(\mu)\otimes\ell^{2}$ with $\|g_{n}-g\|_{\Delta}\to 0$. By assumption (i) we have for each $z\in\Omega$, $\lim_{n\to\infty}bg_{n}(z)=bg(z)$, and that $(g_{n}(z))$ converges weakly to $g(z)$ in $\ell^{2}$, in particular, $\|g(z)\|_{2}\leq\liminf_{n\to\infty}\|g_{n}(z)\|_{2}$. Then since $b(z)$ is a contractive row contraction we can apply Fatou’s lemma to obtain
[…]
This gives
[…]
that is, assumption (ii) is verified as well. Then the result follows by a direct application of Theorem 6. ∎
5.1 Approximation by polynomials in standard weighted sub-Bergman spaces on the unit ball
In this subsection we shall specialize to unitarily invariant weighted Bergman spaces on ${\mathbb{B}}_{d}$, the unit ball in ${\mathbb{C}}^{d}$. These are the spaces ${\mathcal{H}}_{\beta}$ defined in §2.2, where $\beta>d$. Recll that their reproducing kernel is given by
$s^{\beta}(z,w)=\frac{1}{(1-\langle z,w\rangle)^{\beta}},\quad z,w\in{\mathbb{B}}_{d}.$
For such $\beta$ we let $b$ be an analytic row-contraction with $\|b(0)\|<1$. The sub-Bergman space ${{\mathcal{H}}_{\beta}}(b)$ is the Hilbert space ${\mathcal{H}}_{s^{\beta}}(b)$, that is, the Hilbert space with reproducing kernel
$s^{\beta}(z,w)(1-b(z)b(w)^{*})=\frac{1-b(z)b(w)^{*}}{(1-\langle z,w\rangle)^{\beta}}.$
To simplify notations we write ${\mathcal{H}}_{s^{\beta}}(b)={\mathcal{H}}_{\beta}(b)$. Let us note first the following property of the multipliers of these spaces.
Proposition 12.
Let $\beta>d$ and let $b$ an analytic row-contraction with $\|b(0)\|<1$. Then for every $0<\gamma<\beta-d$, $\operatorname{Mult}({\mathcal{H}}_{\gamma})$ is contractively contained in $\operatorname{Mult}({\mathcal{H}}_{\beta}(b))$. In particular, if $u$ is analytic in a neighborhood of $\overline{{\mathbb{B}}_{d}}$ then $u\in\operatorname{Mult}({\mathcal{H}}_{\beta}(b))$.
Proof.
The first assertion is a direct application of Proposition 3 (i). The fact that analytic functions in a neighborhood of $\overline{{\mathbb{B}}_{d}}$ are multipliers of ${\mathcal{H}}_{\gamma},~{}\gamma>0$, is well known and easy to prove. We omit the details. ∎
With this observation in hand we can turn to the approximation result for these spaces. It actually shows that within this context Sarason’s dichotomy does not occur.
Corollary 2.
Let $\beta>d$ and let $b$ an analytic row-contraction with $\|b(0)\|<1$. Then polynomials form a dense subset of ${\mathcal{H}}_{\beta}(b)$. Moreover, for any $e\in\ell^{2}$, the function $z\mapsto b(z)e$ belongs to ${{\mathcal{H}}_{\beta}}(b)$.
Proof.
By Theorem 7, ${\mathcal{H}}_{\beta}(b)$ contains all kernels $s_{w}^{\beta},~{}w\in{\mathbb{B}}_{d}$, of ${\mathcal{H}}_{\beta}$. By Proposition 12 we have $s_{w}^{-\beta}\in\operatorname{Mult}({\mathcal{H}}_{\beta}(b))$, hence $1=s_{w}^{-\beta}s_{w}^{\beta}\in{\mathcal{H}}_{\beta}(b)$. The same argument shows that $1\in{\mathcal{H}}_{\beta-\gamma}(b)$ whenever $\gamma>0,~{}\beta-\gamma>d$. But then Proposition 3 (ii) implies that for such $\gamma$, ${\mathcal{H}}_{\gamma}$ is continuously contained in ${\mathcal{H}}_{\beta}(b)$. Since the kernels $s_{w}^{\beta}$ can be approximated by polynomials in ${\mathcal{H}}_{\gamma}$ the same holds in ${\mathcal{H}}_{\beta}(b)$ and another application of Theorem 7 shows that the closure of polynomials in ${\mathcal{H}}_{\beta}(b)$ contains the dense set $\text{span}\{s_{w}^{\beta}:~{}w\in{\mathbb{B}}_{d}\}$ which concludes the proof.
The previous paragraph shows that $z\mapsto b(z)b(w)^{*}$ belongs to ${{\mathcal{H}}_{\beta}}$ for any $w\in{{\mathbb{B}}_{d}}$. Note that $e\perp b(w)^{*}$ is equivalent to $b(w)e=0$, this proves the last statement of the Corollary. ∎
Using a completely different method, Chu [chu_density] proved the density of polynomials for scalar-valued $b$ in the case when $d=1$ and $\beta=2$.
6 Approximation by polynomials in higher order de Branges-Rovnyak spaces
Let $k$ be a reproducing kernel on a non-empty set ${\mathcal{X}}$, and let $b$ be a contractive multiplier of ${\mathcal{H}}_{k}$. In this section we shall consider kernels of the form
(20) $k^{b,m}(x,y)=k(x,y)(1-b(x)\overline{b(y)})^{m},\quad x,y\in{\mathcal{X}},$
where $m$ is a fixed positive integer. Such kernels have been recently introduced and thoroughly studied in [gu_hypercontractions], in connection with $m-$hypercontractive operators.
Definition 4.
A bounded operator $A$ on the Hilbert space ${\mathcal{H}}$ is called an $m$-hypercontraction if
(21) $\sum_{j=0}^{n}(-1)^{j}\binom{n}{j}A^{*j}A^{j}\geq 0,$
for all $0\leq n\leq m$, where $m,n\in{\mathbb{N}}$.
The result below is a slight simplification of [gu_hypercontractions, Corollary 3.6].
Proposition 13.
Let $b$ be a contractive multiplier of ${\mathcal{H}}_{k}$ such that $|b(x)|<1$ for every $x\in{\mathcal{X}}$. Then $M_{b}^{*}$ is a $m$-hypercontraction if and only if
(22) $k^{b,m}(x,y)=k(x,y)(1-b(x)\overline{b(y)})^{m}\gg 0.$
In this case,
$k(x,y)(1-b(x)b(y))^{n}\gg 0,$
for every integer $n$ with $0\leq n\leq m$.
Proof.
We use repeatedly that the bounded operator $T$ on the reproducing kernel Hilbert space ${\mathcal{H}}_{k}$ is positive if and only if $Tk(x,y)\gg 0$. Thus if $M_{b}^{*}$ is $m-$hypercontractive then (21) holds for $n=m$, hence
$\sum_{j=0}^{m}(-1)^{j}\binom{n}{j}M_{b}^{j}M_{b}^{*j}k(x,y)$ $=\sum_{j=0}^{m}((-1)^{j}\binom{n}{j}b(x)^{j}\overline{b(y)}^{j})k(x,y)$
$=(1-b(x)\overline{b(y)})^{m}k(x,y)\gg 0.$
Conversely, if (22) holds then (21) holds for $n=m$. Moreover, by Schur’s product theorem any positive power of a complete Nevanlinna-Pick kernel is positive, therefore
$\frac{1}{(1-b(x)\overline{b(y)})^{m-n}}\gg 0,$
whenever $n<m$. Then another application of Schur’s theorem gives that the product of the last two positive definite functions is positive definite, i.e.
$k(x,y)(1-b(x)b(y))^{n}\gg 0,$
whenever $n<m$. Equivalently, (21) holds for $0\leq n<m$. Thus $M_{b}^{*}$ is $m-$hypercontractive and the result follows. ∎
We denote by ${\mathcal{H}}_{k}(b,m)$ the Hilbert space with reproducing kernel $k^{b,m}$. We shall investigate polynomial approximation in $k=s^{\beta}$. To simplify notations we denote these spaces by ${\mathcal{H}}_{\beta}(b,m)$. The following result completely solves a conjecture stated in [gu_hypercontractions] in one complex variable ($d=1$).
Corollary 3.
Let $d,m\geq 1$ be integers with $md<\beta$, let $k=s^{\beta}$ and $b\in H^{\infty}({\mathbb{B}}_{d})$ with $\|b\|_{\infty}\leq 1$ and $|b(0)|<1$. Then polynomials form a dense subset of ${\mathcal{H}}_{\beta}(b,m)$.
Proof.
By corollary 2, polynomials form a dense subset in ${\mathcal{H}}_{\beta/m}(b)$, hence by corollary 1 they form a dense subset of ${\mathcal{H}}_{\beta}(b,m)$. ∎
Remark 2.
The result holds when $d=1$ and $m=\beta$, provided that $b$ is a non- extremal point in the unit ball of $H^{\infty}$. Instead of corollary 2 just apply Sarason’s dichotomy [sarason] mentioned in the Introduction.
These methods generalize to cover other situations. For example, letting $b_{1},\ldots,b_{m}$ be non-constant in the unit ball of $H^{\infty}$ one can prove similar results about the space with kernel
$\frac{\prod_{j=1}^{m}(1-b_{j}(z)\overline{b_{j}(w)})}{(1-z\overline{w})^{\beta}},$
where $\beta\geq m$.
Furthermore, we can also consider the case when $b$ is vector-valued. However, in this case, when $\beta=m$ Sarason’s dichotomy is only known to hold when $b$ is a finite dimensional vector.
When $\beta>m$ all results mentioned in this remark extend to the context of several complex variables.
Question 1.
If $d=1$, $\beta=m$ and $b(z)=z$, then clearly ${\mathcal{H}}_{\beta}(b,m)={\mathbb{C}}$. Does there exist $m>1$ and $b$ extremal in the unit ball of $H^{\infty}$ such that ${\mathcal{H}}_{\beta}(b,m)$ contains all polynomials? Do they form a dense subset?
[…]